# Understanding the Complexities of Responsibly Sharing NSFW Content Online

### v1

@@ Abstract
Reddit is in the minority of mainstream social platforms that permit posting content that may be considered to be at the edge of what is ermissible, including so-called Not Safe For Work (NSFW) content. However, NSFW is becoming more common on mainstream platforms, with X now allowing such material. We examine the top 15 NSFW-restricted subreddits by size to explore the complexities of responsibly sharing adult content, aiming to balance ethical and legal considerations with monetization opportunities. We find that users often use NSFW subreddits as a social springboard, redirecting readers to private or specialized adult social platforms such as Telegram, Kik or OnlyFans for further interactions. They also directly negotiate image “trades” through credit cards or payment platforms such as PayPal, Bitcoin or Venmo. Disturbingly, we also find linguistic cues linked to non-consensual content sharing. To help platforms moderate such behavior, we trained a RoBERTa-based classification model, which outperforms GPT-4 and traditional classifiers such as logistic regression and random forest in identifying non-consensual content sharing, demonstrating superior performance in this specific task. The source code and trained model weights are publicly available at https://github.com/socsys/15NSFW˙Subreddits
@@ 1 Introduction
Explicit content is prevalent across the Web — Adult-related websites host user generated explicit content (tyson2016measurements), and some also incorporate social networking features (tyson2015social). This is not illegal in many jurisdictions, although it is common to have age restrictions (e.g., over 18). Since such content can, even if legal, be seen as objectionable by many, the acronym NSFW (Not Safe for Work) is often used as a trigger warning for material that is inappropriate for public or professional environments (cauteruccio2022extraction).
One reason for the huge amount of adult content is its profitability. The online adult entertainment industry was estimated to have grown larger than Netflix, and all of Hollywood in 2018 (guardian), and the growth continues in 2024 as well (yahoo). Regardless of the monetary incentive, mainstream platforms have stayed away from hosting such content. For instance, social networks such as Facebook have policies restricting NSFW content (fbpolicy).
@@
To this end, we turn to Reddit, which occupies a unique position as being a well-known and widely used platform, yet having a highly flexible and permissive policy around content. Each so-called subreddit is allowed to define its own content policy for its user community, which is then (supposed to be) enforced by its moderators. Also Reddit does have a global set of rules (reddit-rules), which supersede subreddit policies and require that the content shared be legal (reddit-rules, Rule 7), consensually posted (reddit-rules, Rule 3) and have appropriate labels for explicit or graphic content (reddit-rules, Rule 6). However, prior work (matias2019civic) highlights the unique emotional and ethical labor involved in moderating sensitive content, with NSFW subreddits often presenting distinct moderation challenges such as content ambiguity, platform scrutiny, and heightened reputational risk.
@@
There are several NSFW subreddits which permit explicit content. We focus on so-called restricted subreddits where only authorised users are allowed to post. In theory, users allowed to post in restricted subreddits would have been vetted by the moderators, allowing us to analyze the complexities of responsibly sharing NSFW content in controlled digital spaces. Analyzing these subreddits provides insight into how a mainstream platform navigates the challenges of managing NSFW content within restrictive environments.
@@ 3. RQ3
We also find posts which directly trade content, using credit cards and payment platforms such as Paypal or Bitcoin, often mentioning explicit \$ values. To promote such posts, users often engage in coordinated campaigns, posting the same content from different accounts, and also posting the same content across different subreddits. Unfortunately, a number of posts appear to promote malicious or illegal content, which could include illegal file-shares, unauthorized adult content etc.  which may violate platform guidelines or local laws.
Finally, we uncover disturbing evidence of a small but nonetheless significant number of posts engaged in what might be non-consensual sharing. We identify such non-consensual posts first from a manual examination and develop RoBERTa-based classifiers (liu1907roberta) to detect and classify such posts. Our model achieves a Recall score of of 0.86. In alignment with the protocol approved by our ethics review panel, we reported these posts to an organization helping Revenge Pornography victims, and upon their advice, we have also informed local law enforcement. Our model is made available freely to help moderators, platforms and regulators detect non-consensual sharing practices at https://github.com/socsys/15NSFW˙Subreddits.
@@ 2 Related Work
Highlighting the need for technology-driven solutions to detect non-consensual content, work by Mohanty et al. (mohanty2019photo) proposes a deep learning-based system for identifying revenge pornography images, combining nudity detection with face recognition to track down such images across platforms. This is complemented by broader research on non-consensual pornography, such as a systematic mapping study (falduti2023mapping) that categorizes technical approaches to combating this issue and identifies gaps in current responses, particularly within cybercrime and digital forensics.
@@ 3.1 Data crawling and pre-processing
3.2 Annotation of non-consensual posts
Given the extensive size of the dataset, manual annotation proved to be labor-intensive. So we selected 5 random subreddits out of the 15 in our dataset for manual annotation: r/curvy, r/Hotwife, r/Naughtywives, WomenBendingOver and r/wifesharing. To streamline the process, we derive keywords (see Table 2) indicating non-consensual posts from manual analysis. Moreover, we observe that users often use intentional misspellings, such as ”le@k” for ”leaks” and ”R4PE” for ”rape,” to evade detection. These modifications are part of leet speak (1337 speak) (leet_wikipedia) that typically involves replacing letters with visually similar numbers or symbols to bypass content moderation. To address this challenge, we normalize the text by converting such symbolic variations back into their standard forms before applying keyword filtering. This step identified 3718 posts, which are then manually annotated by three annotators. Three independent annotators evaluate the posts, labeling them as 0:consensual{0:consensual} or 1:nonconsensual{1:non-consensual}. 298 posts are labeled as non-consensual by at least two annotators. We later use this dataset to train a RoBERTa-based model to detect non-consensual sharing (Section 6.1).
@@ 4.1 Mentions of non-Reddit social handles
Fig. 3(b) illustrates the top 15 social platforms mentioned. An asterisk preceding a platform name denotes a public platform, which is typically used for more visible, wide-reaching interactions. Public platforms like Instagram, X, and Facebook primarily serve to showcase public personas. Private platforms, such as WhatsApp, Telegram, Snapchat, and OnlyFans, cater to a different user intent by providing more controlled, secure environments for one-on-one or small-group interactions, which can potentially be monetized. Kik emerges as the dominant platform, accounting for over 30% of mentions, likely due to its ease of use, emphasis on anonymity, and its suitability for less moderated content.
@@ 4.2 Campaign-like patterns in username sharing
(b) CDF of Gini Coefficients Across Users.
(c) Author distribution per campaign.
(d) Campaign duration
Figure 4: llustrate the distribution, authorship, and temporal dynamics of social media username mentions in Reddit posts. (a) shows that while most usernames are rarely repeated, a small fraction dominate the space. (b) reveals a similar inequality in subreddit participation, with a small user base contributing the majority of activity. (c) suggests that high-frequency usernames are often posted by many different users, hinting at viral or coordinated behavior. (d) supports this further by showing compressed posting time windows for these high-frequency usernames—consistent with campaign-driven or automated posting patterns.
@@
Figure 4(d) presents a Cumulative Distribution Function (CDF) of the time gaps (in days) between posts for having same social media usernames. Tighter Curves for Higher Post Counts, especially [31–50] and >>50, exhibit significantly steeper CDFs. This indicates that most of their posts occur within short time intervals. The purple curve representing usernames mentioned in more than 50 posts rises very quickly, meaning that the vast majority of these posts are made within a narrow time window. This is indicative of coordinated posting behavior, indicating automated or campaign-driven posts. In contrast, usernames mentioned fewer than 5 times (<<5 group) show a broader distribution. Their curve ascends more gradually, suggesting that these mentions are sporadic and distributed over longer time periods—more typical of organic or casual user behavior. The high concentration of posts in short time frames for frequent usernames point to promotional bursts and cross-account coordination, with different Reddit users posting similar content around the same social media username or handle in a synchronized fashion.
@@ 5.2 Prices
We identify 7430 posts that mention the price of content, with prices going up to US\$1,500. Figure 5 presents the top 10 social media platforms mentioned in posts that include price references. OnlyFans dominates with the highest number of price-linked mentions, followed by Snapchat and Kik. Notably, the price ranges and interquartile ranges (IQR) vary widely across platforms, indicating differing monetization strategies or market norms—for instance, WhatsApp shows high pricing variability (IQR: \$134), while OnlyFans exhibits a narrower price spread (IQR: \$3). In the remaining price mentioning posts, users typically either request to be contacted via direct message, share their usernames for other social media platforms, or engage in direct negotiation within the Reddit post itself—for example, posts like ”Would you pay \$80 to see a face?” illustrate explicit pricing discussions. Table 3 presents various payment platforms mentioned across subreddits, along with the number of posts in which each platform is referenced.
@@
Table 4 presents examples of posts referencing content trading. Post such as “Selling Teen Nudes (16–18),” which suggest the potential circulation of sexually explicit material involving minors on NSFW subreddits. Although the age of consent varies across jurisdictions, Reddit’s policies explicitly prohibit any sexually explicit content involving individuals under the age of 18. Threrefore, we reported about the potential presence of sensitive/illegal content to a relevant charity as well as law enforcement authorities.
@@ 6 RQ3: Traces of non-consensual sharing
The rules of Reddit overall, clearly state that consent must be obtained for any content posted (reddit-rules). Yet, our manual analysis identified multiple forms of non-consensual content sharing: 1 Requests for identification: Posts that explicitly ask for help in identifying someone in a photograph or video (“Does anyone know who this is?” or “Help to identify this wife”). 2 Mentions of unawareness: Posts that mention the subject being unaware of the content being shared (“She doesn’t know I’m posting this,” or “got access to her phone”) are strong indicators of non-consensual sharing, highlighting the lack of consent and awareness of the subject, making the content potentially harmful and unethical. 3 Mentions of leaked content: Mentions of “leaked content” often signal the illicit release of private material, typically without the subject’s consent. 4 Mentions of hacking: References to “hacking” in online discourse frequently denote activities that compromise personal accounts or digital privacy, leading to unauthorized access and distribution of content.
@@ 6.1 Classifier training
    Random Forest with TF-IDF representation, implemented using scikit-learn package, and parameters optimized using via GridSearchCV with 5-fold cross validation.
@@ 6.2 Results
Table 5 presents the results. Roberta-based classifiers outperform other classifiers in all evaluation metrics except GPT-4 for Recall. The higher Recall exhibited by GPT-4 indicates that it is more likely to classify negative instances as positive, leading to a greater number of false positives reflected by the lowest value of precision. While a high recall rate can be valuable in ensuring that most relevant instances are detected, it comes with a trade-off in the context of NSFW subreddits. Specifically, this tendency towards false positives means that GPT-4 may incorrectly flag or block posts that actually comply with the subreddit’s guidelines. This could result in unintended censorship of permissible content, undermining the intended purpose of the community and potentially excluding posts that align with the rules of the subreddit. However, this could still be useful with manual oversight, as we discuss below.
@@
Suggested Directions: As NSFW content starts to scale in volume with the new policies on platforms such as X, we believe that classifiers such as the one we have developed will be an essential component in the toolkit of human moderators. We present our initial models only as a baseline for what is possible and advocate further research into identifying better tools and methods to assist moderators in keeping online conversations healthy.
Given the importance of detecting non-consensual practices online, we believe that models with higher recall are more valuable, to ensure the model minimizes missed positive instances. Such a classifier can then be used as a tool that can assist moderators of NSFW platforms. With human oversight, the moderators can be in control of deciding whether or not something is really a false positive, ensuring accurate identification of potential violations. It is important to retain human oversight as completely automated methods could inadvertently censor wrong posts, or miss identifying truly non-consensual ones.
@@ 7 Conclusion
Not Safe For Work (NSFW) content has long been a sensitive category around which there are prohibitions or restrictions on many mainstream social platforms. However, recently there has been a shift in this position, with X relaxing its rules and permitting consensual sharing of NSFW content (Xpolicy). To understand how this might play out, we turn to Reddit, one of the few mainstream platforms which has allowed sharing of NSFW content on clearly marked subreddits. In partricular, we focus on NSFW subreddits which are restricted, allowing only vetted users to share content, which should lead to responsible sharing of NSFW content, including respecting rules around consent.
Interestingly, we found that some of the approved posters on such restricted subreddits are using it as a springboard for drawing users to other platforms such as Only Fans or Kik, where such content can be monetized. Others are using payment mechanisms such as Venmo, Apple Cash, or cryptocurrency to directly trade on the platform, obtaining up to \$1300. Many such users skirting the rules use throwaway accounts, which enables them to rejoin with a different user ID if banned. Throwaway accounts are also used for coordinated advertisement campaigns, violating the sharing spirit of the community.
@@
Discussion: Through this focused study of the largest (restricted) NSFW subreddits, each with a subscriber base of over 1 million users, we aim to contribute to the broader discourse on ethical content sharing and moderation within online communities. Although NSFW subreddits typically intend to share NSFW content freely and consensually, commerce and monetization becomes an important angle, changing the character of the community. Consent, which is an important ground rule on all of Reddit (reddit-rules), appears to have been ignored in a small but significant number of cases. This has obvious implications for other mainstream platforms such as X, which are now starting to allow NSFW content. If NSFW content enters mainstream, it may present new monetization opportunities for platforms, but the impact on the rest of the platform, and its character should be considered. It will become even more important to ensure consent, since it may be easier to detect NSFW content more easily than whether such content was always shared in a legal and consensual manner.

### v2

@@ Abstract
Reddit is in the minority of mainstream social platforms that permit posting content that may be considered to be at the edge of what is permissible, including so-called Not Safe For Work (NSFW) content. However, NSFW is becoming more common on mainstream platforms, with X now allowing such material. We examine the top 15 NSFW-restricted subreddits by size to explore the complexities of responsibly sharing adult content, aiming to balance ethical and legal considerations with monetization opportunities. We find that users often use NSFW subreddits as a social springboard, redirecting readers to private or specialized adult social platforms such as Telegram, Kik or OnlyFans for further interactions. They also directly negotiate image “trades” through credit cards or payment platforms such as PayPal, Bitcoin or Venmo. Disturbingly, we also find linguistic cues linked to non-consensual content sharing. To help platforms moderate such behavior, we trained a RoBERTa-based classification model that outperforms GPT-4 and traditional classifiers, such as logistic regression and random forest, in identifying non-consensual content sharing, showing better performance in this specific task. The source code and model weights are available at https://github.com/socsys/15NSFWsubreddits.
@@ 1 Introduction
Explicit content is prevalent across the Web — Adult-related websites host user generated explicit content (tyson2016measurements), and some also incorporate social networking features (tyson2015social). This is not illegal in many jurisdictions, although it is common to have age restrictions (e.g., 18 years or older). Since such content can, even if legal, be seen as objectionable by many, the acronym NSFW (Not Safe for Work) is often used as a trigger warning for material that is inappropriate for public or professional environments (cauteruccio2022extraction).
One reason for the huge amount of adult content is its profitability. The online adult entertainment industry was estimated to have grown larger than Netflix, and all of Hollywood in 2018 (guardian), and the growth continues in 2024 as well (yahoo). Regardless of the monetary incentive, mainstream platforms have generally avoided hosting such content. For instance, social networks such as Facebook have policies restricting NSFW content (fbpolicy).
@@
To this end, we turn to Reddit, which occupies a unique position as being a well-known and widely used platform, yet having a highly flexible and permissive policy around content. Each so-called subreddit is allowed to define its own content policy for its user community, which is then (supposed to be) enforced by its moderators. Also, Reddit does have a global set of rules (reddit-rules), which supersede subreddit policies and require that the content shared be legal (reddit-rules, Rule 7), consensually posted (reddit-rules, Rule 3) and have appropriate labels for explicit or graphic content (reddit-rules, Rule 6). However, prior work (matias2019civic) highlights the unique emotional and ethical labor involved in moderating sensitive content, with NSFW subreddits often presenting distinct moderation challenges such as content ambiguity, platform scrutiny, and heightened reputational risk.
@@
There are several NSFW subreddits that permit explicit content. We focus on so-called restricted subreddits where only authorised users are allowed to post. In theory, users allowed to post in restricted subreddits would have been vetted by the moderators, allowing us to analyze the complexities of responsibly sharing NSFW content in controlled digital spaces. Analyzing these subreddits provides insight into how a mainstream platform navigates the challenges of managing NSFW content within restrictive environments.
@@ 3. RQ3
We also find posts which directly trade content, using credit cards and payment platforms such as PayPal or Bitcoin, often mentioning explicit \$ values. To promote such posts, users often engage in coordinated campaigns, posting the same content from multiple accounts and also sharing the same content across different subreddits. Unfortunately, a number of posts appear to promote malicious or illegal content, which could include illegal file-shares, unauthorized adult content etc.  which may violate platform guidelines or local laws.
Finally, we uncover disturbing evidence of a small but nonetheless significant number of posts engaged in what might be non-consensual sharing. We identify such non-consensual posts first from a manual examination and develop RoBERTa-based classifiers (liu1907roberta) to detect and classify such posts. Our model achieves a Recall score of of 0.86. In alignment with the protocol approved by our ethics review panel, we reported these posts to an organization helping Revenge Pornography victims, and upon their advice, we have also informed local law enforcement. Our model is made available freely to help moderators, platforms, and regulators detect non-consensual sharing practices at https://github.com/socsys/15NSFWsubreddits.
@@ 2 Related Work
Highlighting the need for technology-driven solutions to detect non-consensual content, work by Mohanty et al. (mohanty2019photo) proposes a deep learning-based system for identifying revenge pornography images, combining nudity detection with face recognition to track down such images across platforms. This is complemented by broader research on non-consensual pornography, such as a systematic mapping study (falduti2023mapping) that categorizes technical approaches to combating this issue and identifies gaps in current responses, particularly within the fields of cybercrime and digital forensics.
@@ 3.1 Data crawling and pre-processing
3.2 Annotation of non-consensual posts
Given the extensive size of the dataset, manual annotation proved to be labor-intensive. So we selected 5 random subreddits out of the 15 in our dataset for manual annotation: r/curvy, r/Hotwife, r/Naughtywives, WomenBendingOver and r/wifesharing. To streamline the process, we derive keywords (see Table 2) indicating non-consensual posts from manual analysis. Moreover, we observe that users often use intentional misspellings, such as ”le@k” for ”leaks” and ”R4PE” for ”rape,” to evade detection. These modifications are part of leet speak (1337 speak) (leet_wikipedia) that typically involves replacing letters with visually similar numbers or symbols to bypass content moderation. To address this challenge, we normalize the text by converting such symbolic variations back into their standard forms before applying keyword filtering. This step identified 3718 posts, which were then manually annotated by three annotators. Three independent annotators evaluate the posts, labeling them as 0:consensual{0:consensual} or 1:nonconsensual{1:non-consensual}. 298 posts are labeled as non-consensual by at least two annotators. We later use this dataset to train a RoBERTa-based model to detect non-consensual sharing (Section 6.1).
(b) CDF of Gini Coefficients Across Users.
(c) Author distribution per campaign.
(d) Campaign duration
Figure 4: Illustrate the distribution, authorship, and temporal dynamics of social media username mentions in Reddit posts. (a) shows that while most usernames are rarely repeated, a small fraction dominates the space. (b) reveals a similar inequality in subreddit participation, with a small user base contributing the majority of activity. (c) suggests that high-frequency usernames are often posted by many different users, hinting at viral or coordinated behavior. (d) supports this further by showing compressed posting time windows for these high-frequency usernames—consistent with campaign-driven or automated posting patterns.
@@ 4.1 Mentions of non-Reddit social handles
Fig. 3(b) illustrates the top 15 social platforms mentioned. An asterisk preceding a platform name denotes a public platform, which is typically used for more visible, wide-reaching interactions. Public platforms like Instagram, X, and Facebook primarily serve to showcase public personas. Private platforms, such as WhatsApp, Telegram, Snapchat, and OnlyFans, cater to different user intents by providing more controlled and secure environments for one-on-one or small-group interactions, which can potentially be monetized. Kik emerges as the dominant platform, accounting for over 30% of mentions, likely due to its ease of use, emphasis on anonymity, and its suitability for less moderated content.
@@ 4.2 Campaign-like patterns in username sharing
Figure 4(d) presents a Cumulative Distribution Function (CDF) of the time gaps (in days) between posts for having the same social media usernames. Tighter Curves for Higher Post Counts, especially [31–50] and >>50, exhibit significantly steeper CDFs. This indicates that most of their posts occur within short time intervals. The purple curve representing usernames mentioned in more than 50 posts rises rapidly, indicating that the vast majority of these posts are made within a narrow time window. This is indicative of coordinated posting behavior, indicating automated or campaign-driven posts. In contrast, usernames mentioned fewer than 5 times (<<5 group) show a broader distribution. Their curve ascends more gradually, suggesting that these mentions are sporadic and distributed over longer time periods—more typical of organic or casual user behavior. The high concentration of posts in short time frames for frequent usernames points to promotional bursts and cross-account coordination, with different Reddit users posting similar content around the same social media username or handle in a synchronized fashion.
@@ 5.2 Prices
We identify 7,430 posts that mention the price of content, with prices going up to US\$1,500. Figure 5 presents the top 10 social media platforms mentioned in posts that include price references. OnlyFans dominates with the highest number of price-linked mentions, followed by Snapchat and Kik. Notably, the price ranges and interquartile ranges (IQR) vary widely across platforms, indicating differing monetization strategies or market norms—for instance, WhatsApp shows high pricing variability (IQR: \$134), while OnlyFans exhibits a narrower price spread (IQR: \$3). In the remaining price-mentioning posts, users typically either request to be contacted via direct message, share their usernames for other social media platforms, or engage in direct negotiation within the Reddit post itself—for example, posts like ”Would you pay \$80 to see a face?” illustrate explicit pricing discussions. Table 3 presents various payment platforms mentioned across subreddits, along with the number of posts in which each platform is referenced.
@@
Table 4 presents examples of posts referencing content trading. Post such as “Selling Teen Nudes (16–18),” which suggest the potential circulation of sexually explicit material involving minors on NSFW subreddits. Although the age of consent varies across jurisdictions, Reddit’s policies explicitly prohibit any sexually explicit content involving individuals under the age of 18. Therefore, we reported about the potential presence of sensitive/illegal content to a relevant charity as well as law enforcement authorities.
@@ 6 RQ3: Traces of non-consensual sharing
The rules of Reddit overall, clearly state that consent must be obtained for any content posted (reddit-rules). Yet, our manual analysis identified multiple forms of non-consensual content sharing:
1
Requests for identification: Posts that explicitly ask for help in identifying someone in a photograph or video (“Does anyone know who this is?” or “Help to identify this wife”). 2 Mentions of unawareness: Posts that mention the subject being unaware of the content being shared (“She doesn’t know I’m posting this,” or “got access to her phone”) are strong indicators of non-consensual sharing, highlighting the lack of consent and awareness of the subject, making the content potentially harmful and unethical. 3 Mentions of leaked content: Mentions of “leaked content” often signal the illicit release of private material, typically without the subject’s consent. 4 Mentions of hacking: References to “hacking” in online discourse frequently denote activities that compromise personal accounts or digital privacy, leading to unauthorized access and distribution of content.
@@ 6.1 Classifier training
    Random Forest with TF-IDF representation, implemented using the scikit-learn package, and parameters optimized using GridSearchCV with 5-fold cross-validation.
@@ 6.2 Results
Table 5 presents the results. Roberta-based classifiers outperform other classifiers in all evaluation metrics except GPT-4 for Recall. The higher Recall exhibited by GPT-4 indicates that it is more likely to classify negative instances as positive, leading to a greater number of false positives reflected by the lowest value of precision. While a high recall rate can be valuable in ensuring that the most relevant instances are detected, it comes with a trade-off in the context of NSFW subreddits. Specifically, this tendency towards false positives means that GPT-4 may incorrectly flag or block posts that actually comply with the subreddit’s guidelines. This could result in unintended censorship of permissible content, undermining the intended purpose of the community and potentially excluding posts that align with the subreddit’s rules. However, this could still be useful with manual oversight, as we discuss below.
@@
Suggested Directions: As NSFW content starts to scale in volume with the new policies on platforms such as X, we believe that classifiers such as the one we have developed will be an essential component in the toolkit of human moderators. We present our initial models only as a baseline for what is possible and advocate further research into identifying better tools and methods to assist moderators in maintaining healthy online conversations.
Given the importance of detecting non-consensual practices online, we believe that models with higher recall are more valuable, to ensure the model minimizes missed positive instances. Such a classifier can then be used as a tool that can assist moderators of NSFW platforms. With human oversight, the moderators can be in control of deciding whether or not something is really a false positive, ensuring accurate identification of potential violations. It is important to retain human oversight as completely automated methods could inadvertently censor wrong posts, or miss-identify truly non-consensual ones.
@@ 7 Conclusion
Not Safe For Work (NSFW) content has long been a sensitive category around which there are prohibitions or restrictions on many mainstream social platforms. However, recently there has been a shift in this position, with X relaxing its rules and permitting consensual sharing of NSFW content (Xpolicy). To understand how this might play out, we turn to Reddit, one of the few mainstream platforms that has allowed sharing of NSFW content on clearly marked subreddits. In particular, we focus on NSFW subreddits which are restricted, allowing only vetted users to share content, which should lead to responsible sharing of NSFW content, including respecting rules around consent.
Interestingly, we found that some of the approved posters on such restricted subreddits are using it as a springboard for drawing users to other platforms such as OnlyFans or Kik, where such content can be monetized. Others are using payment mechanisms such as Venmo, Apple Cash, or cryptocurrency to directly trade on the platform, obtaining up to \$1300. Many such users skirting the rules use throwaway accounts, which enable them to rejoin with a different user ID if banned. Throwaway accounts are also used for coordinated advertisement campaigns, violating the sharing spirit of the community.
@@
Discussion: Through this focused study of the largest (restricted) NSFW subreddits, each with a subscriber base of over 1 million users, we aim to contribute to the broader discourse on ethical content sharing and moderation within online communities. Although NSFW subreddits typically intend to share NSFW content freely and consensually, commerce and monetization become an important angle, changing the character of the community. Consent, which is an important ground rule on all of Reddit (reddit-rules), appears to have been ignored in a small but significant number of cases. This has obvious implications for other mainstream platforms such as X, which are now starting to allow NSFW content. If NSFW content enters mainstream, it may present new monetization opportunities for platforms, but the impact on the rest of the platform and its character should be considered. It will become even more important to ensure consent, since it may be easier to detect NSFW content than to determine whether such content was always shared in a legal and consensual manner.